# Unifying Autoregressive and Diffusion-Based Sequence Generation

### v1

@@ Abstract
We present significant extensions to diffusion-based sequence generation models, blurring the line with autoregressive language models. We introduce hyperschedules, which assign distinct noise schedules to individual token positions, generalizing both autoregressive models (e.g., GPT) and conventional diffusion models (e.g., SEDD, MDLM) as special cases. Second, we propose two hybrid token-wise noising processes that interpolate between absorbing and uniform processes, enabling the model to fix past mistakes, and we introduce a novel inference algorithm that leverages this new feature in a simplified context inspired from MDLM. To support efficient training and inference, we design attention masks compatible with KV-caching. Our methods achieve state-of-the-art perplexity and generate diverse, high-quality sequences across standard benchmarks, suggesting a promising path for autoregressive diffusion-based sequence generation.
@@ 1 Introduction
Generative diffusion models, primarily recognized for their impressive image generation performance in continuous domains (Yang et al., 2023), are rapidly gaining traction in language modeling, a discrete domain historically dominated by autoregressive (AR) models such as the GPT family (Radford et al., 2019; Brown et al., 2020). Contrary to the perceived visceral separation between autoregressive and diffusion models, and despite their distinct historical development, this work reveals a fundamental connection: autoregressive models are, in essence, a form of diffusion.
@@
Despite the common use of Gaussian noise in continuous diffusion, the underlying principles can be adapted to discrete state spaces (Austin et al., 2021; Zhou et al., 2023; Lou et al., 2024). Common practices for sequence generation have the noising process randomly and independently substituting some original tokens by completely unrelated ones, i.e., uniformly sampled tokens or a special “mask” absorbing state. A noise schedule determines token replacement probabilities at different points in the curriculum. The resulting sequence at the schedule’s highest noise level retains no mutual information with the original sequence. The generator is then trained on this curriculum to enable the production of novel sequences.
This work unifies AR and diffusion sequence generation by introducing hyperschedules, allowing different positions in the sequence to be affected by different noise schedules. We establish that autoregressive models, such as GPT, can be understood as diffusion models without data augmentation, utilizing a discrete noise schedule comprising only “full noise” and “no noise” levels. This unification expands model design space and enables a variety of generalized AR-like approaches.
@@
Figure 1: Generative diffusion models prescribe (through $q_{\smash{t|t+1}}$) a curriculum process $\{\mathbf{X}_{t}\}$, then learn (through $p_{\smash{t+1|t}}^{\boldsymbol{\theta}}$) a reverse process $\{\hat{\mathbf{X}}_{t}\}$ so that the marginal distributions match at each step $t$ (vertical squiggly lines). $\mathbf{Y}$ is the training dataset and $\hat{\mathbf{Y}}$ is the generated output. This work focuses on discrete diffusion for sequence generation: our $\mathbf{Y}$, $\hat{\mathbf{Y}}$, $\mathbf{X}_{t}$ and $\hat{\mathbf{X}}_{t}$ are all sequences of discrete tokens, using the identity for both $q_{T}$ and $\pi$. We show that standard autoregressive models (e.g., GPT) are an extreme case of this framework, a unification enabling a vast continuum of new diffusion models, including autoregressive-like ones.
Recent (Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and concurrent (Liu et al., 2024; Kim et al., 2025; Peng et al., 2025; Nie et al., 2025; Wang et al., 2025; Arriola et al., 2025) works have focused on mask diffusion models (MDMs). By specializing on the “absorb” noising mechanism, these MDMs enable great simplifications over a general-purpose treatment: neural networks no longer need an explicit noise-level dependency, and a more standard loss function can be used. However, our work shows that the same feats and simplifications can be accomplished in a general, non-MDM case. Motivated by the same rationale that led concurrent MDMs to conceive “remasking” strategies, we introduce hybrid noising processes, interpolating between the “absorb” and “uniform” processes to combine the benefits of both and achieve state-of-the-art performances, with aspects further improved by our novel adaptive correction sampler (ACS) inference algorithm. Our hyperschedule-equipped approach also supports specialized attention matrices enabling KV-caching and efficient training.
@@
$\boldsymbol{\tau}_{\text{Quench}}$
(a) Quenched (AR).
$\boldsymbol{\tau}_{\text{Flat}}$
(b) Flat annealing.
$\boldsymbol{\tau}_{\text{Block}}^{\omega{=4}}$
(c) Block annealing.
$\boldsymbol{\tau}_{\text{Slide}}^{\omega{=4}}$
(d) Slide annealing.
Figure 2: We introduce $\boldsymbol{\tau}$-hyperschedules, subjecting different token positions $i$ with different noise levels (red high; blue low) at different generation step $t$. (2(a)) Standard AR models (e.g., GPT) determine tokens one by one, “quenching” each of them to full determination in a single step. They may thus be construed as an extreme case of a diffusion model. (2(b)) Standard diffusion models (e.g., SEDD) gradually anneal all tokens independently of their position. (2(c)) Block-wise application of flat annealing, here for blocks of width $\omega=4$. (2(d)) Annealing with a sliding window (“smoothed” AR), here using window width $\omega=4$. These last two examples share important features of both AR and diffusion models.
2 Abstract Sequence Generators
@@
In this section, we reconcile autoregressive and diffusion-based sequence generation models by abstracting-out their respective implementation details, instead emphasizing their shared essence. Note that we focus on unconditional generation without loss of generality; conditional generation (e.g., prompting) may be recovered as a special case.
2.1 Generative Diffusion Models
For our present purpose, a generator is a procedure that yields an output $\hat{\mathbf{y}}$ according to a certain probability distribution $\operatorname{\mathds{P}}_{\hat{\mathbf{Y}}}$. We focus on procedures composed of $T\in\mathds{N}^{*}$ discrete steps (e.g., “calls” to a neural network) each updating a state $\hat{\mathbf{x}}_{t}$ to a state $\hat{\mathbf{x}}_{t+1}$ using a conditional probability distribution $\operatorname{\mathds{P}}_{\hat{\mathbf{X}}_{t+1}|\hat{\mathbf{X}}_{t}}(\hat{\mathbf{x}}_{t+1}|\hat{\mathbf{x}}_{t})=p_{t+1|t}^{\boldsymbol{\theta}}(\hat{\mathbf{x}}_{t+1}|\hat{\mathbf{x}}_{t})$ with learned parameters $\boldsymbol{\theta}$. The stochastic process $\{\hat{\mathbf{X}}_{t}\}$ is seeded with an initial state $\hat{\mathbf{x}}_{0}$ that is either a constant or sampled from a provided (i.e., not learned) distribution $\operatorname{\mathds{P}}_{\mathbf{X}_{0}}(\mathbf{x}_{0})=p_{0}(\mathbf{x}_{0})$. The output $\hat{y}=\pi(\hat{x}_{T})$ is a deterministic function of $\hat{x}_{T}$. Given a dataset sampled from a data distribution $\operatorname{\mathds{P}}_{\mathbf{Y}}$, our goal is to train the parameters $\boldsymbol{\theta}$ so that the generator’s marginal output distribution $\operatorname{\mathds{P}}_{\hat{\mathbf{Y}}}$ matches the data’s $\operatorname{\mathds{P}}_{\mathbf{Y}}$ (hereafter noted ${{\hat{\mathbf{Y}}}}$${{\mathbf{Y}}}$).
In general, generative diffusion models Yang et al. (2023) prescribe the evolution $\operatorname{\mathds{P}}_{\mathbf{X}_{t}|\mathbf{X}_{t+1}}(\mathbf{x}_{t}|\mathbf{x}_{t+1})=q_{t|t+1}(\mathbf{x}_{t}|\mathbf{x}_{t+1})$ of a curriculum $\{\mathbf{X}_{t}\}$ conditional on a data sample $\mathbf{y}$ according to $\operatorname{\mathds{P}}_{\mathbf{X}_{T}|\mathbf{Y}}(\mathbf{x}_{T}|\mathbf{y})=q_{T}(\mathbf{x}_{T}|\mathbf{y})$. [^1] Note that $t$ here goes down from $T$ to 00 as we align our notation with generation steps. Existing diffusion work often take the opposite perspective, tracking a “noise level” $T-t$. Different training strategies have been developed to align the marginal distributions ${{\hat{\mathbf{X}}_{t}}}$${{\mathbf{X}_{t}}}$ at every step $t$. In a sense, the curriculum $\{\mathbf{X}_{t}\}$ breaks the goal ${{\hat{\mathbf{Y}}}}$${{\mathbf{Y}}}$ into $T$ simpler steps. Figure 1 provides a diagrammatic summary.
2.2 Sequence Generators
This work focuses on datasets of sequences $\mathbf{y}=(y^{0},y^{1},\cdots,y^{d-1})\in\mathcal{Y}^{d}$ composed of $d\in\mathds{N}^{*}$ tokens from a finite set $\mathcal{Y}$. Although the curriculum $\mathbf{x}_{t}=(x_{t}^{0},\cdots,x_{t}^{d-1})\in\mathcal{X}^{d}$ may in general rely on a continuous $\mathcal{X}$, all our explicit examples use $\mathcal{X}=\mathcal{Y}$, $q_{T}(\mathbf{y}|\mathbf{y})=1$ and $\pi(\hat{\mathbf{x}}_{T})=\hat{\mathbf{x}}_{T}$. In any case, the output sequence $\hat{\mathbf{y}}\in\mathcal{Y}^{d}$ and states $\mathbf{x}_{t}\in\mathcal{X}^{d}$ match these domains. We further focus on curriculum-prescribing processes that factorize in terms of per-token transitions $q_{t|t+1}^{i}(x^{\prime}|x)$, i.e., $q_{t|t+1}(\mathbf{x}_{t}|\mathbf{x}_{t+1})=\prod_{i}q_{t|t+1}^{i}(x_{t}^{i}|x_{t+1}^{i})$. We now consider two important examples.
Standard AR.
Most modern language models predict tokens autoregressively: one token at a time, each conditional on the tokens that precede it. This is an extreme case of a diffusion model where the conditional probability $q_{t|t+1}$ is actually a deterministic function that masks-out the $t$-th token, ultimately resulting in $\mathbf{x}_{0}$ composed solely of MASK tokens. At generation, $\hat{\mathbf{x}}_{0}$ starts as all masks, and each $p_{t+1|t}^{\boldsymbol{\theta}}(\hat{\mathbf{x}}_{t+1}|\hat{\mathbf{x}}_{t})$ predicts the sole entry $\hat{x}_{t+1}^{t}$ that differs between $\hat{\mathbf{x}}_{t+1}$ and $\hat{\mathbf{x}}_{t}$. Implementations need not actually track MASK tokens because they always end the sequence. Training on a sample $\mathbf{y}$ uses cross-entropy loss on $\operatorname{\mathds{P}}_{\hat{X}_{i+1}^{i}|\hat{\mathbf{X}}_{i}^{:i}}(y^{i}|\mathbf{y}^{:i})$ for each position $i$ [^2] Indices a:b go from $a$ (inclusive) to $b$ (exclusive); omitted $a$ or $b$ are implicit 00 or $d$, respectively.
Standard diffusion.
@@ 2.2 Sequence Generators
Discrete diffusion models prescribes the curriculum using $q_{t|t+1}^{i}(x^{\prime}|x)$ given by the $x$-th column of a matrix $Q_{t|t+1}$ that is independent of the position $i$. For example, SEDD (Lou et al., 2024) uses $Q_{t|t+1}=\exp\bigl{(}(\bar{\sigma}_{T-t}-\bar{\sigma}_{T-t-1})Q_{\text{tok}}\bigr{)}$, where the transition matrix $Q_{\text{tok}}$ is one of two possibilities
@@
and $\bar{\sigma}_{0}\leq\bar{\sigma}_{1}\leq\cdots\leq\bar{\sigma}_{T}$ are cumulative noise schedules such that $\bar{\sigma}_{0}\approx 0$ and $\bar{\sigma}_{T}\approx\infty$. Here $\lvert\mathcal{Y}\rvert$ is the number of tokens in the set $\mathcal{Y}$, including the special MASK token associated with the last dimension of $Q_{\text{tok}}$. In words, using $Q_{\text{Uniform}}$ gradually replaces tokens by random ones, while $Q_{\text{Absorb}}$ gradually replaces them by MASK. Generation starts from $\hat{\mathbf{x}}_{0}$ sampled from the stationary distribution $p_{0}$ (i.e., random non-mask tokens for $Q_{\text{Uniform}}$ and all-masks for $Q_{\text{Absorb}}$), and $p_{t+1|t}^{\boldsymbol{\theta}}$ is learned in terms of diffusion weighted denoising score entropy (DWDSE) (Lou et al., 2024).
@@
2.3 Transformers
Like most modern language models, all sequence generators considered in this work are implemented as transformers (Vaswani et al., 2017), more specifically the backbone from Peebles & Xie (2023). We call Aligned the transformer configuration used in most masked language models and diffusion models: each transformer cell predicts (output) the same token position as the one it receives (input). Conversely, we call Shifted the configuration used in most autoregressive models: each cell predicts the next token in the sequence. To simplify discussions, we fully commit to our sequence generator abstraction when indexing positions, irrespective of these input/output configuration (see details in Appendix B).
@@ 3 Autoregressive Sequence Diffusion
(a) $d=8$ tokens using $\boldsymbol{\tau}_{\text{Block}}^{\omega{=4}}$.
(b) $d=6$ tokens using $\boldsymbol{\tau}_{\text{Slide}}^{\omega{=4}}$.
Figure 3: Examples of our hyperschedule-powered hybrid model in action. Gold tokens are settled; dark grey mask tokens are worthless; and the remaining tokens are active. Active mask (light gray) are “known unknowns”. Here the “uniform” component of our hybrid process introduces one error in the curriculum (red), teaching the model to fix such errors at inference. Although both hyperschedules generate at the limit rate of $\rho=1$ token per step in the long-sequence limit, $\boldsymbol{\tau}_{\text{Slide}}^{\omega{=4}}$ experiences an initial overhead of $\omega-1$ steps.
3.1 Hyperschedules
We generalize standard diffusion curricula by subjecting different token positions $i$ to different noise schedules, distinguishing the number of generation steps $T$ from the number of noise levels $\mathcal{T}$. Concretely, $q_{t|t+1}^{i}(x^{\prime}|x)$ is given by the $x$-th column of a $Q_{t|t+1}^{i}$ obtained by substituting each instance of $\sigma_{T-t}$ or $\alpha_{T-t}$ in $Q_{t|t+1}$ by $\sigma\raisebox{2.0pt}{${}_{\tau_{t}^{i}}$}$ or $\alpha\raisebox{2.0pt}{${}_{\tau_{t}^{i}}$}$, respectively, where the hyperschedule $\tau_{t}^{i}\in\{0,1,\cdots,\mathcal{T}\}$ satisfying $\mathcal{T}=\tau_{0}^{i}\geq\tau_{1}^{i}\geq\cdots\geq\tau_{T}^{i}=0$ for all positions $i\in\{0,\cdots,d-1\}$. In effect, the noise schedule ($\bar{\boldsymbol{\sigma}}$ or $\boldsymbol{\alpha}$) unfolds differently at different positions. Figure 2 provides some examples.
We introduce two characterizations of an hyperschedule. First, we define the window width $\omega$ as the largest (among all steps $t$) number of positions $i$ for which $(\tau_{t}^{i},\tau_{t+1}^{i})$ is neither $(0,0)$ nor $(\mathcal{T},\mathcal{T})$. All other things being equal, a lower $\omega$ offers more opportunities to improve inference time (see Sec. 3.3 for examples). Standard AR models use $\boldsymbol{\tau}_{\text{Quench}}$ with value $1$ where $i\geq t$ and 00 elsewhere, and thus have $\omega=1$ by construction (Fig. 2(a)). Standard diffusion models use $\boldsymbol{\tau}_{\text{Flat}}$ with value $\mathcal{T}-t$ (often called “noise level” or “time”; not to be confused with our generation step $t$) for all $i$, using $\mathcal{T}=T$ and $\omega=d$. Two more examples are provided, both parametrized by $\omega$: concurrent work on block diffusion (Arriola et al., 2025) may be understood in terms of $\boldsymbol{\tau}_{\text{Block}}^{\omega{}}$ (Fig. 2(c)), and we introduce novel $\boldsymbol{\tau}_{\text{Slide}}^{\omega{}}$ (Fig. 2(d)).

### v2

@@ Abstract
We present significant extensions to diffusion-based sequence generation models, blurring the line with autoregressive language models. We introduce hyperschedules, which assign distinct noise schedules to individual token positions, generalizing both autoregressive models (e.g., GPT) and conventional diffusion models (e.g., SEDD, MDLM) as special cases. Second, we propose two hybrid token-wise noising processes that interpolate between absorbing and uniform processes, enabling the model to fix past mistakes, and we introduce a novel inference algorithm that leverages this new feature in a simplified context inspired from MDLM. To support efficient training and inference, we design attention masks compatible with KV-caching. Our methods achieve state-of-the-art perplexity and generate diverse, high-quality sequences across standard benchmarks, suggesting a promising path for autoregressive diffusion-based sequence generation. See code and resources at https://hdlm-colm.github.io/.
@@ 1 Introduction
Generative diffusion models, primarily recognized for their impressive image generation performance in continuous domains (Yang et al., 2023), are rapidly gaining traction in language modeling, a discrete domain historically dominated by autoregressive (AR) models such as the GPT family (Radford et al., 2019; Brown et al., 2020). Contrary to the apparent separation between AR and diffusion models, and despite their distinct historical development, this work reveals a fundamental connection: AR models are a form of diffusion.
@@
Despite the common use of Gaussian noise in continuous diffusion, the underlying principles can be adapted to discrete state spaces (Austin et al., 2021; Zhou et al., 2023; Lou et al., 2024). Common practices for sequence generation have the noising process randomly and independently substituting some original tokens by completely unrelated ones, i.e., uniformly sampled tokens or a special “mask” absorbing state. A noise schedule determines token replacement probabilities at different points in the curriculum. The resulting sequence at the schedule’s highest noise level retains no mutual information with the original sequence. The generator is trained on this curriculum to enable the production of novel sequences.
This work unifies AR and diffusion sequence generation by introducing hyperschedules, allowing different positions in the sequence to be affected by different noise schedules. We establish that autoregressive models, such as GPT, can be understood as diffusion models without data augmentation, utilizing a discrete noise schedule comprising only “full noise” and “no noise” levels. This unification expands model design space and enables a variety of generalized AR-like approaches. With the hindsight, hyperschedule-like concepts can be identified in the “Mask” process of Bansal et al. (2023) as well as in the “Swin-DPM” technique of Feng et al. (2024).
@@
Figure 1: Generative diffusion models prescribe (through $q_{\smash{t|t+1}}$) a curriculum process $\{\mathbf{X}_{t}\}$, then learn (through $p_{\smash{t+1|t}}^{\boldsymbol{\theta}}$) a reverse process $\{\hat{\mathbf{X}}_{t}\}$ so that the marginal distributions match at each step $t$ (vertical squiggly lines). $\mathbf{X}_{T}$ is the training dataset and $\hat{\mathbf{X}}_{T}$ is the generated output. This work focuses on discrete diffusion for sequences of discrete tokens. We show that standard autoregressive models (e.g., GPT) are an extreme case of this framework, a unification enabling a vast continuum of diffusion models, including autoregressive ones.
Recent (Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and concurrent (Liu et al., 2024; Kim et al., 2025; Peng et al., 2025; Nie et al., 2025; Wang et al., 2025; Arriola et al., 2025) works have focused on mask diffusion models (MDMs). By specializing on the “absorb” noising mechanism, these MDMs enable great simplifications over a general-purpose treatment: neural networks no longer need an explicit noise-level dependency, and a more standard loss function can be used. However, our work shows that the same feats and simplifications can be accomplished in a general, non-MDM case. Motivated by the same rationale that led concurrent MDMs to conceive “remasking” strategies, we introduce hybrid noising processes, interpolating between the “absorb” and “uniform” processes to combine the benefits of both and achieve state-of-the-art performances, with aspects further improved by our novel adaptive correction sampler (ACS) inference algorithm. Our hyperschedule-equipped approach supports specialized attention matrices enabling KV-caching and efficient training.
@@
Figure 2: We introduce $\boldsymbol{\tau}$-hyperschedules (top row) – subjecting different token positions $i$ with different noise levels (red high; blue low) at different generation step $t$ – and illustrate how three different noising processes (bottom 3 rows) can be modulated by such hyperschedules. Hyperschedules. (a) Standard AR models (e.g., GPT) determine tokens one by one, “quenching” each of them to full determination in a single step. They may thus be construed as an extreme case of a diffusion model. (b) Standard diffusion models (e.g., SEDD) gradually anneal all tokens independently of their position. (c) Block-wise application of flat annealing, here for blocks of width $\omega=3$. (d) Annealing with a sliding window (“smoothed” AR), here using window width $\omega=3$. These last two examples share important features of both AR and diffusion models. While the 4 presented examples all generate $\rho=1$ token per step in the long-sequence limit (with the caveat that slide experiences an initial overhead of $\omega-1$ steps), the last 3 patterns are all straightforwardly adapted to $\rho>1$ (“quick draft”) and $\rho<1$ (“think hard”) regimes. Noising Processes. (2) The absorb noising process – a.k.a. Masked Diffusion Model (MDM) – overwrites tokens with a special MASK token. These masks are “known unknowns”: it is clear to a denoising model that it must put a non-mask token in its stead. Conversely, unmasked tokens are taken as “absolute truth” during generation: once a token has been unmasked, it remains unaltered until the end. (3) The uniform noising process overwrites with (non-mask) tokens selected uniformly at random from the vocabulary of possible tokens. The orange/light-blue color coding is not part of the state $\mathbf{x}_{t}$, and is provided solely for the reader’s convenience. The model thus has no direct way to know if a token has been altered before (“unknown unknowns”), and may thus revisit a position’s value many times during generation. (4) The hybrid noising process blends a little bit of uniform into the absorb process. When denoising, MASK tokens still represent clear “known unknowns”, while unmasked ones have become “a priori good” candidates that may however need to be “fixed”. We argue that it is desirable for models to learn to fix their own mistakes.
2 Unification Through Abstraction
@@
In this section, we reconcile autoregressive and diffusion-based language models by abstracting-out their respective implementation details, instead emphasizing their shared essence.
Sequence Generator
We define a sequence generator as a stochastic procedure that repeatedly calls a neural network to yield a sequence of $d$ tokens, each taken from a finite vocabulary $\mathcal{X}$. We write $\hat{\mathbf{x}}_{t}=(\hat{x}_{t}^{0},\cdots,\hat{x}_{t}^{d-1})\in\mathcal{X}^{d}$ the state of the generator after $t$ calls: starting from a (trivial) initial state $\hat{\mathbf{x}}_{0}$, the neural network specifies the probability $p_{t+1|t}^{\boldsymbol{\theta}}(\hat{\mathbf{x}}_{t+1}|\hat{\mathbf{x}}_{t})$ to get to state $\hat{\mathbf{x}}_{t+1}$ from state $\hat{\mathbf{x}}_{t}$. The generated sequence $\hat{\mathbf{x}}_{T}$ is obtained after $T$ such calls. The bottom row of Figure 1 illustrates this process (from right to left, writing $\hat{\mathbf{X}}_{t}$ the stochastic variable taking value $\hat{\mathbf{x}}_{t}$).
Discrete Diffusion Sequence Generators
A sequence generator’s neural network must learn the function $p_{t+1|t}^{\boldsymbol{\theta}}(\hat{\mathbf{x}}_{t+1}|\hat{\mathbf{x}}_{t})$ such that the yielded $\hat{\mathbf{x}}_{T}$ approximates the distribution of clean training samples $\mathbf{x}_{T}$ (no hat!) from the training dataset. To this end, the “diffusion” paradigm (Yang et al., 2023) is to prescribe a distribution $q_{t-1|t}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ from which we can sample $\mathbf{x}_{t-1}$ given $\mathbf{x}_{t}$. This noising process (top row in Figure 1) gradually destroys the information in the training data sample $\mathbf{x}_{T}$, until nothing remains of it in $\mathbf{x}_{0}$; we call curriculum the process $\{\mathbf{X}_{t}\}$ obtained by recursively sampling the prescribed $q_{t-1|t}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ starting from a training sample $\mathbf{x}_{T}$. The training objective for $p_{\smash{t+1|t}}^{\boldsymbol{\theta}}$ amounts to denoising such noisy samples, aiming to align the marginal distribution of $\hat{\mathbf{X}}_{t}$ with its $\mathbf{X}_{t}$ counterpart (noted $\hat{\mathbf{X}}_{t}$$\mathbf{X}_{t}$}} in Figure 1).
Token-Wise Processes
@@
The noising process is typically constructed out of many independent processes, each affecting one of the original tokens, i.e., $q_{t|t+1}(\mathbf{x}_{t}|\mathbf{x}_{t+1})=\prod_{i}q_{t|t+1}^{i}(x_{t}^{i}|x_{t+1}^{i})$. In practice, $q_{t|t+1}^{i}(x_{t}^{i}|x_{t+1}^{i})$ is given by the $x$-th column of a matrix $Q_{t|t+1}$ that is independent of the position $i$. For example, SEDD (Lou et al., 2024) uses $Q_{t|t+1}=\exp\bigl((\bar{\sigma}_{T-t}-\bar{\sigma}_{T-t-1})Q_{\text{tok}}\bigr)$, where the transition matrix $Q_{\text{tok}}$ is either of
@@
and $\bar{\sigma}_{0}\leq\bar{\sigma}_{1}\leq\cdots\leq\bar{\sigma}_{T}$ are cumulative noise schedules such that $\bar{\sigma}_{0}\approx 0$ and $\bar{\sigma}_{T}\approx\infty$. Here $\lvert\mathcal{X}\rvert$ is the number of tokens in the set $\mathcal{X}$, including a special MASK token associated with the last dimension of $Q_{\text{tok}}$. In words, using $Q_{\text{Uniform}}$ gradually replaces tokens by random ones, while $Q_{\text{Absorb}}$ gradually replaces them by MASK. Generation starts from $\hat{\mathbf{x}}_{0}$ sampled from the stationary distribution $p_{0}$ (i.e., random non-mask tokens for $Q_{\text{Uniform}}$ and all-masks for $Q_{\text{Absorb}}$), and $p_{t+1|t}^{\boldsymbol{\theta}}$ is learned in terms of diffusion weighted denoising score entropy (DWDSE) (Lou et al., 2024).
@@
AR Sequence Generators
Most modern language models predict tokens autoregressively: one token at a time, each conditional on the tokens that precede it. This may be viewed as a sequence generator with $T=d$: $\hat{\mathbf{x}}_{0}=(\text{MASK},\text{MASK},\cdots,\text{MASK})$ represents an empty sequences using a special MASK token, and each $p_{t+1|t}^{\boldsymbol{\theta}}(\hat{\mathbf{x}}_{t+1}|\hat{\mathbf{x}}_{t})$ predicts the token $\hat{x}_{t+1}^{t}$ to substitute the first MASK encountered in $\hat{\mathbf{x}}_{t}$. This can be viewed as an extreme case of a diffusion model where the conditional probability $q_{t|t+1}$ is actually a deterministic function that masks-out the $t$-th token, ultimately resulting in $\mathbf{x}_{0}$ composed solely of MASK tokens. However, notice that this $q_{t|t+1}(\mathbf{x}_{t}|\mathbf{x}_{t+1})=\prod_{i}q_{t|t+1}^{i}(x_{t}^{i}|x_{t+1}^{i})$ involves factors depending on the position $i$, which is forbidden in the discussion above; Section 3.1 palliates to this issue.
Transformers
All sequence generators considered in this work are implemented as transformers (Vaswani et al., 2017), more specifically the backbone from Peebles & Xie (2023). One subtlety is that many diffusion and/or masked language models use an Aligned configuration where each transformer cell predicts (output) the same token position as the one it receives (input). In contrast, AR models use the Shifted configuration: each cell predicts the next token in the sequence. These subtleties can be abstracted out at an high level (see details in Figure 4 in Appendix).
@@ 3 Autoregressive Sequence Diffusion
3.1 Hyperschedules: position-dependent schedules
We generalize standard diffusion curricula by subjecting different token positions $i$ to different noise schedules, distinguishing the number of generation steps $T$ from the number of noise levels $\mathcal{T}$. Concretely, $q_{t|t+1}^{i}(x^{\prime}|x)$ is given by the $x$-th column of a $Q_{t|t+1}^{i}$ obtained by substituting each instance of $\sigma_{T-t}$ or $\alpha_{T-t}$ in $Q_{t|t+1}$ by $\sigma\raisebox{2.0pt}{${}_{\tau_{t}^{i}}$}$ or $\alpha\raisebox{2.0pt}{${}_{\tau_{t}^{i}}$}$, respectively, where the hyperschedule $\tau_{t}^{i}\in\{0,1,\cdots,\mathcal{T}\}$ satisfying $\mathcal{T}=\tau_{0}^{i}\geq\tau_{1}^{i}\geq\cdots\geq\tau_{T}^{i}=0$ for all positions $i\in\{0,\cdots,d-1\}$. In effect, the noise schedule ($\bar{\boldsymbol{\sigma}}$ or $\boldsymbol{\alpha}$) unfolds differently at different positions; hyperschedules act as schedules for schedules. See examples on top row of Fig. 2.
We introduce two characterizations of an hyperschedule. First, we define the window width $\omega$ as the largest (among all steps $t$) number of positions $i$ for which $(\tau_{t}^{i},\tau_{t+1}^{i})$ is neither $(0,0)$ nor $(\mathcal{T},\mathcal{T})$. All other things being equal, a lower $\omega$ offers more opportunities to improve inference time (see Sec. 3.3 for examples). Standard AR models use $\boldsymbol{\tau}_{\text{Quench}}$ with value $1$ where $i\geq t$ and 0 elsewhere, and thus have $\omega=1$ by construction (Fig. 2(a.1)). Standard diffusion models use $\boldsymbol{\tau}_{\text{Flat}}$ with value $\mathcal{T}-t$ (often called “noise level”) for all $i$, using $\mathcal{T}=T$ and $\omega=d$ (Fig. 2(b.1)). Two $\omega$-parametrized examples are also provided: concurrent work on block diffusion (Arriola et al., 2025) may be understood in terms of $\boldsymbol{\tau}_{\text{Block}}^{\omega{}}$ (Fig. 2(c.1)), and we introduce novel $\boldsymbol{\tau}_{\text{Slide}}^{\omega{}}$ (Fig. 2(d.1)).